# Multi-head Spatial-Spectral Mamba for Hyperspectral Image Classification

### v1

@@ Abstract
I Introduction
The extensive spectral data in Hyperspectral Images (HSIs) presents both challenges and opportunities for effective classification [1]. The successful application of Neural Networks has bolstered recent progress in HSIC [2, 3, 4, 5, 6], with a growing interest in Transformer models to enhance HSI analysis. The Transformer architecture has significantly advanced HSI Classification (HSIC) compared to traditional deep learning (TDL) methods [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Transformers, with their self-attention mechanisms, can capture long-range dependencies within spatial-spectral features, enabling a comprehensive understanding of the intricate relationships between spectral bands and spatial information [18]. The quadratic computational complexity of Transformers presents significant challenges, particularly when dealing with high-dimensional HSI data. This complexity can limit their practical applications. Additionally, Transformers typically require a large number of labeled data for training to achieve high performance; otherwise, they are prone to overfitting.
The emergence of Mamba, a state space model-based (SSM) approach, offers a promising solution to the challenges of high-dimensional HSI data. Mamba can capture long-range dependencies like Transformers but with significantly greater computational efficiency, making them well-suited for processing large datasets without compromising performance [19, 20, 21]. Originally designed for sequence data in natural language processing, Mamba models have proven effective by incorporating time-varying parameters into SSMs, achieving linear time complexity while maintaining strong representational capabilities [22]. This innovative approach has been extended to various visual tasks, including HSIC, where TDL often struggles with the data’s high dimensionality and complexity [23].
Figure 1: A joint spatial-spectral feature token is first computed from the hyperspectral image (HSI). These tokens are then encoded in the MHSSMamba model, which includes token enhancement and a multi-head attention module, allowing for a more selective and effective representation of information compared to standard fixed-dimension encodings. The output is subsequently processed through a state-space model, followed by normalization and a linear layer, before being passed to the classification head for ground truth generation.
@@ I Introduction
Given the limitations of the Mamba architecture and the complexities of HSI data, this work introduces the Spatial-Spectral Mamba for HSIC. This new architecture builds on the original Mamba framework, addressing its shortcomings by incorporating token enhancement and multi-head self-attention mechanisms. The key contributions includes:
  • The Spatial-Spectral Mamba with Multi-head Self-Attention (MHSSMamba) enhances spectral tokens, improving feature representation and classification performance.
  • MHSSMamba balances spectral and spatial features by generating and integrating both types of tokens using multi-head attention, capturing complex relationships between spectral bands and spatial locations.
  • MHSSMamba reduces overfitting through token generation, which lowers input dimensionality, and a feature enhancement mechanism that focuses on important features, ensuring robustness even with limited training data.
  • MHSSMamba also captures long-range dependencies within HSI data, improving pattern recognition across different image regions. Additionally, state-space modeling effectively handles the sequential nature of HSI data, maintaining contextual information across spectral bands for better feature extraction and classification.
@@
II Spatial-Spectral Mamba with Multi-Head Self-Attention
@@
Assume the HSI data has a shape of $(H,W,C)$, where $H$ and $W$ are the height and width, and $C$ is the number of bands. Table I and Figure 1 provide an overview of the MHSSMamba method. The HSI cube $X$ is divided into overlapping 3D patches as:
$N=\bigg{(}\frac{H}{P}\times\frac{W}{P}\bigg{)}$ (1)
where $Patch(X)\in\mathcal{R}^{N\times(P\times P\times C)}$. These patches are divided into spectral and spatial patches, each undergoing separate processing to generate spectral and spatial tokens.
TABLE I: Summary of the Mathematical Expressions
Variable Details
Input HSI Cube $\textbf{X}\in\mathcal{R}^{B\times H\times W\times C}$
Spatial Tokens: $\mathbf{S}$ $[\mathbf{s}_{1},\mathbf{s}_{2},...,\mathbf{s}_{C}]\in\mathbb{R}^{B\times(HW)\times C}$
Spectral Token: $\mathbf{F}$ $[\mathbf{f}_{1},\mathbf{f}_{2},...,\mathbf{f}_{HW}]\in\mathbb{R}^{B\times(HW)\times C}$
Spatial Enhancement $\widetilde{\mathbf{S}}^{(l)}$ $\mathbf{S}^{(l)}\odot\sigma(\mathbf{W}_{s}\mathbf{c}+\mathbf{b}_{s})$
Spectral Enhancement $\widetilde{\mathbf{F}}^{(l)}$ $\mathbf{F}^{(l)}\odot\sigma(\mathbf{W}_{f}\mathbf{c}+\mathbf{b}_{f})$
Multi-Head Attention $O$ $Concat(O_{1},O_{2},\dots,O_{h})$
State Transition $h_{t}$ $ReLU(W_{transition}h_{t-1}+W_{update}E_{t})$
Final Output $y$ $\sigma(h_{t}W_{classifier})$
II-A Spectral-Spatial Token Generation:
@@
HSIs have rich spatial and spectral information. By flattening 2D spatial data into spatial tokens $\mathbf{S}$ and 1D spectral data into spectral tokens $\mathbf{F}$, the model captures intricate spatial-spectral relationships. This token representation enables the multi-head attention module to jointly attend to spatial and spectral cues, which is crucial for accurate HSI classification. We generate spatial and spectral tokens $\mathbf{S}$ and $\mathbf{F}$ as:
$\mathbf{S}=[\mathbf{s}_{1},\mathbf{s}_{2},...,\mathbf{s}_{C}]\in\mathbb{R}^{B\times(HW)\times C}$ (2)
$\mathbf{F}=[\mathbf{f}_{1},\mathbf{f}_{2},...,\mathbf{f}_{HW}]\in\mathbb{R}^{B\times(HW)\times C}$ (3)
II-B Token Enhancement and Multi-head Attention
This module applies a gating mechanism to spatial and spectral tokens, using the center region of the HSI sample as context [19]. This allows MHSSMamba to adjust the importance of different tokens dynamically, enhancing feature extraction. Given the center tokens, the token enhancements are made as follows:
@@ II-B Token Enhancement and Multi-head Attention
where $W_{s}$ and $W_{f}$ are learned weights and $\sigma$ is the sigmoid function. The multi-head attention mechanism enables the model to learn diverse and informative feature representations by attending to different parts of the input tokens (both spatial and spectral). By projecting the input tokens into multiple attention heads, the model captures complex dependencies within HSI data, essential for effective feature extraction. The scaled dot-product attention, followed by a softmax operation, dynamically assigns importance weights to the input tokens, enhancing the model’s focus on the most relevant features. Let the spatial and spectral tokens $\widetilde{\mathbf{S}}^{(l)}$ and $\widetilde{\mathbf{F}}^{(l)}$ be the inputs, and $Q,K,V$ be the query, key, and value respectively. For each head $i$: $Q_{i}=\widetilde{\mathbf{S}}^{(l)}W^{Q}_{i}$, $K_{i}=\widetilde{\mathbf{F}}^{(l)}W^{K}_{i}$, and $V_{i}=\widetilde{\mathbf{F}}^{(l)}W^{V}_{i}$, where $W^{Q}_{i},W^{K}_{i},W^{V}_{i}$ are learned weight matrices. The attention scores are computed as:
@@
where the attention output is $O_{i}=A_{i}V_{i}$ while the concatenated heads are $O=Concat(O_{1},O_{2},\dots,O_{h})$.
II-C State Space Model
Given a sequence of enhanced tokens $O=(E_{1},E_{2},E_{3},\dots,E_{T})$, the state transition is computed as:
@@ II-C State Space Model
where $W_{transition}$ and $W_{update}$ are learned weights, and $h_{t}$ is the hidden state at time $t$. The final output is obtained by applying a linear classifier on $h_{t}$:
@@
III Experimental Results and Discussion
The proposed MHSSMamba model is evaluated using several publicly available Hyperspectral datasets including The WHU-Hi-LongKou [28, 29], The University of Pavia (UP), The Salinas (SA), and The University of Houston (UH) datasets. MHSSMamba’s weights were initialized randomly and optimized over 50 epochs using the Adam optimizer with a learning rate of 0.001 and softmax loss. The training was done in mini-batches of 256 samples per epoch. The Mamba block’s embedding dimensions were set to 64, with 4 heads for multi-head attention and state space dimensions of 128. This setup allowed MHSSMamba to learn patterns by adjusting its parameters to minimize loss. Training, validation, and test samples were randomly selected, with various random split percentages tested. The initial patch size was $4\times 4$, but different sizes were also evaluated, and an optimized set of samples and patch sizes was chosen for comparative results. All experiments were conducted on Google Colab, utilizing a Python 3 notebook with a GPU, 25 GB of RAM, and 358.27 GB of cold storage. Figure 2 shows the convergence loss and accuracy of our proposed MHSSMamba model over 50 epochs.
Figure 2: The accuracy and loss for both the training and validation sets were computed using a $6\times 6$ patch and 25% training samples over 50 epochs on the University of Houston dataset.
Table II presents MHSSMamba’s performance with different training and test percentages using a $4\times 4$ patch size and various patch sizes. Larger patches capture minute details and local patterns but are prone to noise and overfitting, especially with smaller samples. Smaller patches encapsulate global features and contextual information, enhancing robustness against noise. As sample sizes increase, MHSSMamba’s robustness and generalization improve. Computational performance is influenced by internet speed and available RAM. Figures 3, 4, 5, and 6 show qualitative results, with quantitative results in Table II.
TABLE II: The performance of the proposed model is evaluated using various patch sizes and different test splits. For all test splits, a $4\times 4$ patch size is used over 50 epochs with a batch size of 256. For all patch sizes, 10% of the data is allocated for training and validation samples.
Tr/Va/Te(%) AA OA $\kappa$ Time (s) Patch AA OA $\kappa$ Time (s)
@@ III-A Comparative Methods
TABLE III: University of Houston Classification Results over several SOTA methods.
@@
TABLE IV: Pavia University Classification Results over several SOTA methods.
@@
IV Conclusions

### v3

@@ Abstract
articletype: Letter
1 Introduction
Hyperspectral Image Classification (HSIC) is essential in various fields due to its capacity to capture detailed spectral information across numerous narrow bands, enabling precise material identification and analysis. This capability has facilitated significant advancements and applications in domains such as remote sensing [1], Earth observation [2], urban planning [3], agriculture [4], mineral exploration [5], environmental monitoring [6], and climate change [7]. Additionally, HSIC is valuable in less geoscience-related fields such as food processing [8, 9], bakery products [10], bloodstain identification [11, 12], and meat processing [13, 14].
The extensive spectral data in Hyperspectral Images (HSIs) presents both challenges and opportunities for effective classification [15]. The successful application of Neural Networks has bolstered recent progress in HSIC [16, 17, 18, 19, 20], with a growing interest in Transformer models to enhance HSI analysis. The Transformer architecture has significantly advanced HSIC compared to traditional deep learning (TDL) methods [21, 22, 23, 24, 25, 26, 27, 28, 29]. Transformers, with their self-attention mechanisms, can capture long-range dependencies within spatial-spectral features, enabling a comprehensive understanding of the intricate relationships between spectral bands and spatial information. The quadratic computational complexity of Transformers presents significant challenges, particularly when dealing with high-dimensional HSI data. This complexity can limit their practical applications. Additionally, Transformers typically require a large number of labeled data for training to achieve high performance; otherwise, they are prone to overfitting.
The emergence of Mamba, a state space model-based (SSM) approach, offers a promising solution to the challenges of high-dimensional HSI data. Mamba can capture long-range dependencies like Transformers but with significantly greater computational efficiency, making them well-suited for processing large datasets without compromising performance [30, 31]. Originally designed for sequence data in natural language processing, Mamba models have proven effective by incorporating time-varying parameters into SSMs, achieving linear time complexity while maintaining strong representational capabilities [32]. This innovative approach has been extended to various visual tasks, including HSIC, where TDL often struggles with the data’s high dimensionality and complexity [33].
@@ 1 Introduction
In response to the challenges posed by traditional Mamba architecture and the complexities of HSI data, we introduce a multi-head Spatial-Spectral Mamba (MHSSMamba) for HSIC. By building on the Mamba framework, the proposed architecture incorporates cutting-edge token enhancement and multi-head self-attention mechanisms. The MHSSMamba dramatically enhances spectral token representation, leading to superior feature extraction and classification performance. Our key contributions are:
  1. 1.
    The Spectral-Spatial token extracts separate spectral and spatial tokens from the input HSI patches, allowing the MHSSMamba model to independently leverage both types of information. This approach potentially enhances feature representation and improves classification performance.
  2. 2.
    The customized Multi-head self-attention (MHSA) mechanism is specifically designed to process spatial-spectral tokens by projecting queries, keys, and values through dense layers and appropriately reshaping them. This design enables the MHSSMamba to perform attention operations more efficiently and effectively, capturing complex relationships between different spectral bands and spatial locations.
  3. 3.
    The spectral-spatial feature enhancement module introduces a dual gating mechanism that separately enhances spectral and spatial tokens processed by MHSA using learned gating signals. This dual-gate approach enables the model to adaptively refine features based on both spatial and spectral contexts, potentially increasing the discriminative power of the learned representations.
  4. 4.
    The State Space Model (SSM) introduces a novel way of capturing temporal dynamics by maintaining and updating state representations through learned transitions and updates. This model component integrates sequential dependencies into the HSI processing pipeline, enhancing the ability to model temporal patterns in time-series HSI data.
In summary, the MHSSMamba presents a comprehensive end-to-end pipeline that integrates token generation, multi-head attention, feature enhancement, and state space modeling. This hybrid architecture combines several advanced techniques into a unified framework, offering an intricate approach to HSIC that leverages both spatial and spectral information as well as temporal dynamics. These contributions underscore the model’s innovative techniques for addressing HSIC challenges, providing potential improvements in feature representation, attention mechanisms, and temporal modeling.
@@
2 Spatial-Spectral Mamba with Multi-Head Self-Attention
@@
Assume the HSI data has a shape of $(H,W,C)$, where $H$ and $W$ are the height and width, and $C$ is the number of bands. Figure 1 provides an overview of the MHSSMamba method. The HSI cube $X$ is divided into overlapping 3D patches as $N=\big{(}\frac{H}{P}\times\frac{W}{P}\big{)}$, where $Patch(X)\in\mathcal{R}^{N\times(P\times P\times C)}$, $P$ be the size of patch. These patches are divided into spectral and spatial patches, each undergoing separate processing to generate spectral and spatial tokens.
Figure 1: A joint spatial-spectral feature token is first computed from the HSI. These tokens are then encoded in the MHSSMamba model, which includes token enhancement and a multi-head attention module, allowing for a more selective and effective representation of information compared to standard fixed-dimension encodings. The output is subsequently processed through a state-space model, followed by normalization and a linear layer, before being passed to the classification head for ground truth generation.
@@
HSIs have rich spatial and spectral information. By flattening 2D spatial data into spatial tokens $\mathbf{S}$ and 1D spectral data into spectral tokens $\mathbf{F}$, the model captures intricate spatial-spectral relationships. This token representation enables the multi-head attention module to jointly attend to spatial and spectral cues, which is crucial for accurate HSI classification. We generate spatial and spectral tokens $\mathbf{S}$ and $\mathbf{F}$ as $\mathbf{S}=[\mathbf{s}_{1},\mathbf{s}_{2},...,\mathbf{s}_{C}]\in\mathbb{R}^{B\times(HW)\times C}$ and $\mathbf{F}=[\mathbf{f}_{1},\mathbf{f}_{2},...,\mathbf{f}_{HW}]\in\mathbb{R}^{B\times(HW)\times C}$. The token enhancement applies a gating mechanism to spatial and spectral tokens, using the center region of the HSI sample as context [30]. This allows MHSSMamba to adjust the importance of different tokens dynamically, enhancing feature extraction. Given the center tokens, the token enhancements are made as follows:
@@
where $\mathbf{S}^{(l)}$ and $\mathbf{F}^{(l)}$ represents the spatial and spectral tokens at layer $l$ of the model. Spatial and spectral tokens capture spatial and spectral features from the HSI samples, respectively. $\mathbf{c}$ denotes the center region of the HSI sample used as context. It helps to provide additional information about the central part of the patch to guide the gating mechanism. $\mathbf{W}_{s}$ and $\mathbf{W}_{f}$ are the learned weight parameters used to transform the center context $\mathbf{c}$ for spatial and spectral tokens, respectively. It is a matrix that projects the context into the same space as the spatial and spectral tokens. $\mathbf{b}_{s}$ and $\mathbf{b}_{f}$ represent the bias terms added to the transformed context for spatial and spectral tokens. $\sigma$ denotes the sigmoid function, which is used to apply a gating mechanism. The sigmoid function outputs values between 0 and 1, allowing the model to adjust the importance of different tokens dynamically. $\odot$ denotes element-wise multiplication. It is used to apply the gating values (obtained from the sigmoid function) to the spatial and spectral tokens.
The multi-head attention mechanism enables the model to learn diverse and informative feature representations by attending to different parts of the input tokens (both spatial and spectral). By projecting the input tokens into multiple attention heads, the model captures complex dependencies within HSI data, essential for effective feature extraction. The scaled dot-product attention, followed by a softmax operation, dynamically assigns importance weights to the input tokens, enhancing the model’s focus on the most relevant features. Let the spatial and spectral tokens $\widetilde{\mathbf{S}}^{(l)}$ and $\widetilde{\mathbf{F}}^{(l)}$ be the inputs, and $Q,K,V$ be the query, key, and value respectively. For each head $i$: $Q_{i}=\widetilde{\mathbf{S}}^{(l)}W^{Q}_{i}$, $K_{i}=\widetilde{\mathbf{F}}^{(l)}W^{K}_{i}$, and $V_{i}=\widetilde{\mathbf{F}}^{(l)}W^{V}_{i}$, where $W^{Q}_{i},W^{K}_{i},W^{V}_{i}$ are learned weight matrices. The attention scores are computed as:
@@
where the attention output is $O_{i}=A_{i}V_{i}$ while the concatenated heads are $O=Concat(O_{1},O_{2},\dots,O_{h})$. Given a sequence of enhanced tokens $O=(E_{1},E_{2},E_{3},\dots,E_{T})$, the state transition is computed as:
@@
where $O=(E_{1},E_{2},E_{3},\dots,E_{T})$ denotes a sequence of enhanced tokens, where $E_{t}$ represents the token at the time step $t$, and $T$ is the total number of tokens in the sequence. $h_{t}$ represents the hidden state at the time step $t$. It captures the context and dependencies from previous tokens and is updated based on the enhanced token $E_{t}$. $W_{transition}$ is a learned weight matrix used in the transition function. It transforms the previous hidden state $h_{t-1}$ to combine with the current token $E_{t}$ in the state update process. $W_{update}$ is a learned weight matrix applied to the current enhanced token $E_{t}$ during the state update process. $ReLU$ denotes the Rectified Linear Unit activation function, which introduces non-linearity into the model by outputting the maximum of 0 and the input value. It is applied to the weighted sum of the previous hidden state and the current token. The final output is obtained by applying a linear classifier on $h_{t}$:
@@
where $W_{classifier}$ is a learned weight matrix used in the linear classifier. It projects the hidden state $h_{t}$ to obtain the final output. $\sigma$ denotes the sigmoid function, which is used to generate probabilities for classification. It outputs values between 0 and 1. $y$ represents the final output of the model, which is obtained by applying the linear classifier to the hidden state $h_{t}$ and passing it through the sigmoid function.
@@
3 Experimental Results and Discussion
The proposed MHSSMamba model is evaluated using several publicly available Hyperspectral datasets including The WHU-Hi-LongKou [38, 39], The University of Pavia (UP), The Salinas (SA), and The University of Houston (UH) datasets. MHSSMamba’s weights were initialized randomly and optimized over 50 epochs using the Adam optimizer with a learning rate of 0.001 and softmax loss. The training was done in mini-batches of 256 samples per epoch. The Mamba block’s embedding dimensions were set to 64, with 4 heads for multi-head attention and state space dimensions of 128.
This setup allowed MHSSMamba to learn patterns by adjusting its parameters to minimize loss. Training, validation, and test samples were randomly selected, with various random split percentages tested. The initial patch size was $4\times 4$, but different sizes were also evaluated, and an optimized set of samples and patch sizes was chosen for comparative results. All experiments were conducted on Google Colab, utilizing a Python 3 notebook with a GPU, 25 GB of RAM, and 358.27 GB of cold storage.
Table 1 presents MHSSMamba’s performance with different training and test percentages using a $4\times 4$ patch size and various patch sizes. Larger patches capture minute details and local patterns but are prone to noise and overfitting, especially with smaller samples. Smaller patches encapsulate global features and contextual information, enhancing robustness against noise. As sample sizes increase, MHSSMamba’s robustness and generalization improve. Computational performance is influenced by internet speed and available RAM. Figures 2, 3, 4, and 5 show qualitative results, with quantitative results in Table 1.
Table 1: The performance of the proposed model is evaluated using various patch sizes and different test splits. For all test splits, a $4\times 4$ patch size is used over 50 epochs with a batch size of 256. For all patch sizes, 10% of the data is allocated for training and validation samples.
Tr/Va/Te(%) AA OA $\kappa$ Train Time (s) Patch AA OA $\kappa$ Train Time (s)
@@ 3.1 Comparative Methods
Table 2: University of Houston Classification Results over several SOTA methods.
@@
Table 3: Pavia University Classification Results over several SOTA methods.
@@
4 Computational Complexity
The computational complexity of the proposed model can be analyzed by evaluating each of its primary components, for instance, the token generation layer, which generates the spatial-spectral tokens using the dense layer, has a complexity of $O(B\times H\times W\times C\times out\_Channels)$. The Multi-head self-attention layer involves matrix multiplications and attention score calculations, resulting in a complexity of $O(B\times L\times embed\_Dim^{2}+B\times num\_heads\times L^{2}\times head\_dim)$. The feature enhancement layer responsible for enhancing features using dense layers and element-wise operations, contributed $O(B\times L\times out\_Channels^{2})$. Finally, the SSM layer which updates the state through dense layers over $T$ timesteps has a complexity of $O(T\times B\times State\_dim^{2})$. Combining all layers, the overall complexity of the MHSSMamba is $O(B\times H\times W\times C\times out\_Channels^{2}+B\times L\times embed_{d}im^{2}+B\times num\_heads\times L^{2}\times head\_dim+B\times L\times out\_Channels^{2}+T\times B\times state\_dim^{2})$, reflecting the cumulative computational demands of token generation, attention mechanisms, feature enhancement, and state updates. Here we also discuss the best, average, and worst-case scenarios of the proposed model.
Best Case: When the model operates with minimal features and token sizes, and there is no need for extensive processing, the complexity of the token generation is $O(B\times H\times W\times C)$, where $B$ is the batch size, $H,~{}W$ are the spatial dimensions, and $C$ is the number of bands. The multi-head self-attention under the minimal head and embedding sizes has a complexity of $O(B\times L\times embed\_dim)$, and the SSm assuming minimal state dimension and timesteps length operates with $(T\times B\times state\_dim)$. Average Case: With the typical sizes for token generation, the complexity will be $O(B\times H\times W\times C\times out\_Channels)$. the multi-head self-attention considering the standard attention heads and embeddings sizes, results in $O(B\times L\times embed\_dim^{2}+B\times num\_heads\times L^{2}\times head\_dim)$. For the feature enhancement module, the complexity is $O(B\times L\times out\_channels^{2})$, while the SSm operates at $O(T\times B\times state\_dim^{2})$. Worst Case: In scenarios with maximal feature and token sizes, the complexity of the token generation can reach $O(B\times H\times W\times C\times out\_Channels)$. The multi-head self-attention complexity in the worst case is $O(B\times L^{2}\times emded\_dim^{2})$ due to the large attention scores and weights. The feature enhancement layers scale to $O(B\times L\times out\_Channels^{2})$, and the SSM could require $O(T\times B\times state\_dim^{2})$ in the most extensive cases.
5 Conclusions
@@
Disclosure statement
The authors declare no conflict of interest exists.
Data Availability
The data used in this study is publicly available and can be accessed from the corresponding data repository Hyperspectral Datasets.